OpenStudy (anonymous):

how many different groups of 5 players can be chosen from a team of 13?

OpenStudy (mr.math):

\[=\left(\begin{matrix}13 \\ 5\end{matrix}\right)=1287\]

OpenStudy (mr.math):

Generally, the number of size-k subsets of a size-n set is given by \[\left(\begin{matrix}n \\ k\end{matrix}\right)={n! \over k!(n-k)!}\]